# Instructions for *ACL Proceedings

### v1

@@ Abstract
With the widespread adoption of Large Language Models (LLMs) across various applications, it is empirical to ensure their fairness across all user communities. However, most LLMs are trained and evaluated on Western centric data, with little attention paid to low-resource languages and regional contexts. To address this gap, we introduce PakBBQ, a culturally and regionally adapted extension of the original Bias Benchmark for Question Answering (BBQ) dataset. PakBBQ comprises over 214 templates, 17180 QA pairs across 8 categories in both English and Urdu, covering eight bias dimensions including age, disability, appearance, gender, socio-economic status, religious, regional affiliation, and language formality that are relevant in Pakistan. We evaluate multiple multilingual LLMs under both ambiguous and explicitly disambiguated contexts, as well as negative versus non negative question framings. Our experiments reveal (i) an average accuracy gain of 12% with disambiguation, (ii) consistently stronger counter bias behaviors in Urdu than in English, and (iii) marked framing effects that reduce stereotypical responses when questions are posed negatively. These findings highlight the importance of contextualized benchmarks and simple prompt engineering strategies for bias mitigation in low resource settings.
@@ 1 Introduction
Most existing bias benchmarks and fairness evaluations for question answering (QA) systems such as the Bias Benchmark for QA (BBQ)Parrish et al. (2022) have been developed with Western, primarily English speaking contexts in mind. While these resources have been instrumental in revealing cultural and demographic biases, they do not adequately capture the unique social divisions, linguistic nuances, and historical power dynamics present in other regions. As a result, models deployed in low-resource or non-Western settings can exhibit untested and potentially more severe biases toward locally salient groups, such as caste, sect, or clan affiliations as supported in the following works Khandelwal et al. (2023), Ferrara (2023).
@@
Although there have been attempts to contextualize the original BBQ dataset in relation to the local context, little to no work has been done on a QA dataset tailored to the Pakistani context. KO-BBQ Jin et al. (2024) is a culturally adapted Korean version of the BBQ dataset, and is rooted in Korean culture, while its Chinese adapted version CBBQ Huang and Xiong (2024) captures nuances embedded within the Chinese culture. These datasets are not transferrable to Pakistani contexts, specifically due to the diverse social, cultural and language landscape of Pakistan. Pakistan’s rich cultural diversity stems from its multi-ethnic population spread across different provinces, each with distinct languages, traditions, views and socio-economic status as shown in several studies Shah and Amjad (2011) . With regional languages like Punjabi, Sindhi, Pashto, Balochi, and others coexisting alongside Urdu as the national language, and with deep rooted regional identities and cross regional biases, a one-size-fits-all dataset fails to capture the nuanced realities of the country. The effect is further propagated by Pakistan’s religious landscape, dominated by Islam which is split among sects like Barelvi, Deobandi, Christians, Hindus, and other minorities. This sectarian and interfaith diversity creates varied social norms and complicates uniform data representation. As Yaqin (2022) mentions, Urdu encodes social hierarchies through pronouns, honorifics, and register choices (Persian-Arabic vs. Hindustani vocabulary), signaling respect, status, and group membership. Gendered verb and adjective agreement further embeds masculine authority and feminine marginality. Modeling these formality markers in PakBBQ reveals how LLMs may reproduce structural biases along urban–rural, educational, and gender lines unique to Pakistan.
@@
To bridge this gap, we introduce PakBBQ (Pakistani Bias Benchmark for Question Answering), a culturally and regionally adapted bias benchmark for QA tailored to the Pakistani context. Building upon the original BBQ dataset, we imply a methodology similar to the one in KOBBQ to contextualize and adapt the dataset to Pakistani norms and culture. Templates were categorized into: Target Modified (TM) templates adapted for Pakistani context (e.g., replacing Western names with local counterparts), Sample Removed (SR) templates inapplicable locally, and Newly Added (NA) templates capturing Pakistan specific biases (caste, sect, clan, regional affiliations), validated by native speakers. We also remove template categories irrelevant to Pakistani context, and add new categories such as Regional and Language Formality Biases, identified through large scale scrapping of various media articles, research papers, social blogs.
@@
  • PakBBQ Dataset: A collection of 217 templates instantiated into 17180 English and Urdu scripted QA pairs covering 8 bias dimensions specific to Pakistan.
@@
By releasing PakBBQ, covering eight bias dimensions (Age, Disability Status, Language Formality, Gender Identity, Physical Appearance, Regional, Religion, and Socioeconomic Status(SES)), we aim to enable more rigorous auditing and mitigation of social biases in QA models deployed in Pakistan, and to provide a blueprint for culturally sensitive bias benchmarks in other underrepresented regions. The dataset and code are available at PakBBQ.
@@ 2.1 Bias Benchmarks in QA and Cross-Cultural Adaptations
Natural language processing models have been shown to inherit and even amplify societal biases present in their training data, which can manifest in question answering (QA) tasks as stereotypical or discriminatory outputs. Parrish et al. Parrish et al. (2022) introduced the Bias Benchmark for QA (BBQ) to evaluate such biases across nine social dimensions in U.S. English under both under informative and fully informative contexts. Subsequent frameworks, such as UnQover Li et al. (2020), employ underspecified questions to surface biases like gendered name–occupation associations, while pronoun based methods Zhao et al. (2018) reveal gender bias via pronoun usage though these are less applicable in Urdu, which conveys gender through verb and adjective agreement rather than explicit pronouns.
@@ 3 Dataset Construction
3.1 Adaptation Strategy: DT, TM, and NA Categories
@@
Directly Translated (DT): Items in this category were translated into Urdu (or another relevant local language) without significant changes, as their social and cultural contexts were already applicable to Pakistani society. These include examples with globally common scenarios like age-based assumptions or gender stereotypes.
@@
Newly Added (NA): This category includes examples specifically constructed for the Pakistani sociocultural landscape. These involve biases unique to Pakistan, such as sectarian affiliations, regional or ethnic identities (e.g., Sindhi, Baloch), and minority religious groups (e.g., Ahmadis, Hindus). We also incorporate formality biases specific to the Urdu language. Since direct English prompts are not feasible, we use Roman Urdu to evaluate English responses. These examples aim to capture context-specific stereotypes not present in the original BBQ dataset.
@@ 3.2 Template Annotation
For the Newly Added (NA) templates, we employed a structured annotation process involving multiple annotators (undergraduate Pakistani university students) to ensure both cultural relevance and consistency in identifying bias. Annotators were first briefed on the aims of the study and made explicitly aware of the potential risks of exposure to stereotypes, sexism, and other harmful biases contained within the templates.
@@ 3.3 Translation
For the translation of our dataset, we essentially translated the context, question and answer choices. Three translation models were evaluated, Facebook’s seamless0m4t-v2-large Barrault et al. (2023), Gemma3 27B  [^3] https://huggingface.co/google/gemma-3-27b-it Team et al. (2025) and Google Translate API  [^4] https://cloud.google.com/translate. Facebook and Gemma3 models frequently produced semantically distorted outputs, for example incorrectly translating “dark-skinned” as “andhera" ie: "darkness” instead of “Kaala" ie: “dark-skinned”, and failed to preserve contextual cues across Q/As. After extensive manual comparison and evaluation, Google Translate API produced the most accurate and contextually sound translations, thus was chosen to translate our English version of the dataset to Urdu .
@@ 4 Experimental Setup
In the following section, we describe the methodology applied to evaluate the dataset across both languages.
@@ 4.2 Model Selection
We selected a diverse set of latest multilingual LLM models, capable of handling zero-shot question answering. We also ensures that a diversity was maintained in terms of model sizes as well, and the final models used were, DeepSeek-V3(671B)  [^5] https://deepseekv3.org , GPT 4.1 Nano  [^6] https://platform.openai.com/docs/models/gpt-4.1-nano , GPT 4.1 Mini [^7] https://platform.openai.com/docs/models/gpt-4.1-mini, GPT 4.1888https://platform.openai.com/docs/models/gpt-4.1, Gemini 2.0 Flash  [^9] https://deepmind.google/technologies/gemini/flash/ and Gemini 2.0 Flash Lite [^10] https://deepmind.google/technologies/gemini/flash-lite/. While the exact parameters of some of these models have not been disclosed, we ensured that our evaluation covered a representative range of model scales, ranging from lightweight, midsized and large scale LLMs.
@@ 6.1 Accuracy Comparison
ENG GPT4.1-Nano 0.80 0.83 0.68 0.81
ENG GPT4.1-Mini 0.82 0.87 0.62 0.87
ENG GPT4.1 0.82 0.88 0.49 0.89
ENG DeekSeekv3 0.85 0.91 0.55 0.92
ENG gemini-2.0-flash-lite 0.88 0.93 0.61 0.94
ENG gemini-2.0-flash 0.84 0.90 0.57 0.87
UR GPT4.1-Nano 0.72 0.73 0.67 0.74
UR GPT4.1Mini 0.75 0.78 0.61 0.81
UR GPT4.1 0.75 0.78 0.62 0.81
UR Deepseekv3 0.67 0.67 0.50 0.85
UR gemini-2.0-flash-lite 0.69 0.71 0.51 0.77
UR gemini-2.0-flash 0.81 0.85 0.61 0.88
Table 1: Accuracy comparison of various models across English and Urdu
@@ 6.2 Ambig vs Disambig Accuracy
All models gain substantially from disambiguation. Accuracy jumps by about 12% points on average. Urdu models show higher variance than English ($\sigma\approx 0.11$ vs. 0.07), highlighting their sensitivity to prompt clarity. The results on a whole, underscore the value of explicit disambiguation, particularly in lower resource languages.
6.3 Neg vs Non-Neg Accuracy
@@
ENG GPT4.1-Nano 0.83 0.78
ENG GPT4.1-Mini 0.83 0.82
ENG GPT4.1 0.81 0.81
ENG DeekSeekv3 0.85 0.84
ENG gemini-2.0-flash-lite 0.89 0.87
ENG gemini-2.0-flash 0.85 0.82
UR GPT4.1-Nano 0.73 0.71
UR GPT4.1Mini 0.77 0.74
UR GPT4.1 0.77 0.74
UR Deepseekv3 0.69 0.64
UR gemini-2.0-flash-lite 0.71 0.66
UR gemini-2.0-flash 0.81 0.81
Table 2: Comparison of Models repsonding to Negative and Non-Negative Questions
@@ 6.4 Bias Scores Across Categories
Accuracy is not enough to evaluate biases in LLMs as we must also evaluate how biased or counter-biased the model is if it does not choose the Unknown option. We apply the bias score metric used in BBQParrish et al. (2022). The provided heatmaps in Figure  4 illustrate bias scores for models evaluated on both English and Urdu across several bias categories, including Age, Disability Status, Gender Identity, Physical Appearance, Regional Bias, Religion, and Socioeconomic Status (SES), under both ambiguous and disambiguated contexts. The bias score (1-1 to $1$) measures a model’s preference between biased and counter-biased choices (excluding “Unknown”): a score of 0 indicates no preference, values near $1$ indicate bias, and values near 1-1 indicate counter-bias.
@@ 7.4 Disambiguation as a Bias Mitigation Strategy
Explicit disambiguation yields an average accuracy gain of 12 pp across all models (e.g., GPT4.1-Mini in Urdu rises from 64 % to 87 %). This suggests that much of the bias in ambiguous prompts stems from models’ reliance on learned probabilistic defaults rather than deliberate reasoning, and that clear contextual cues enable them to override these assumptions. In practice, disambiguation offers a straightforward prompt-engineering technique, especially valuable for lower-resource languages like Urdu, though it may be less feasible when explicit information is unavailable. The stronger effect observed in Urdu also points to language-specific bias mechanisms, warranting further investigation into how linguistic structure and culture shape model behavior.
@@ 8 Conclusion
This paper introduces PakBBQ, the first culturally adapted bias benchmark for evaluating large language models in the Pakistani context. Our findings reveal substantial performance disparities between English and Urdu (7-17pp accuracy gap) and demonstrate that bias evaluation cannot rely on simple translation of Western frameworks. Pakistan specific bias categories showed consistently poor performance, highlighting the necessity of culturally grounded evaluation. However, we identify practical mitigation strategies, explicit disambiguation improves accuracy by 12pp on average, while negative question framing reduces stereotypical responses—both effects being stronger in Urdu than English. These findings challenge the transferability assumption in AI bias evaluation and provide immediate prompt engineering solutions for more equitable multilingual AI deployment. Our work establishes a replicable methodology for developing culturally specific bias benchmarks, emphasizing the urgent need to move beyond English centric evaluation frameworks toward inclusive approaches that address the diverse realities of global AI systems.
@@
While PakBBQ represents a first step towards culturally grounded bias evalution for Pakistani QA systems, several limitation should be noted. Our dataset is primarily limited to Urdu and English scripts and does not cater towards major regional languages (eg., Punjabi, Sindhi, Pashto, Balochi), which may exhibit distinct bias patterns. Secondly our evaluation adopts a zero-shot prompting strategy with a fixed “You are a Pakistani person” system prompt and relies on the assumption that all models interpret formality and honorific cues consistently in both languages; in practice, morphological and register mismatches may introduce noise. Furthermore, errors and context misalignment is possible during the translation of the dataset, translation errors can arise especially for context sensitive terms like skin tone or sectarian identifiers which may affect bias measurements downstream.

### v2

# PakBBQ: A Culturally Adapted Bias Benchmark for QA

@@ Abstract
With the widespread adoption of Large Language Models (LLMs) across various applications, it is imperative to ensure their fairness across all user communities. However, most LLMs are trained and evaluated on Western centric data, with little attention paid to low-resource languages and regional contexts. To address this gap, we introduce PakBBQ, a culturally and regionally adapted extension of the original Bias Benchmark for Question Answering (BBQ) dataset. PakBBQ comprises over 214 templates, 17180 QA pairs across 8 categories in both English and Urdu, covering eight bias dimensions including age, disability, appearance, gender, socio-economic status, religious, regional affiliation, and language formality that are relevant in Pakistan. We evaluate multiple multilingual LLMs under both ambiguous and explicitly disambiguated contexts, as well as negative versus non negative question framings. Our experiments reveal (i) an average accuracy gain of 12% with disambiguation, (ii) consistently stronger counter bias behaviors in Urdu than in English, and (iii) marked framing effects that reduce stereotypical responses when questions are posed negatively. These findings highlight the importance of contextualized benchmarks and simple prompt engineering strategies for bias mitigation in low resource settings.
@@ 1 Introduction
Most existing bias benchmarks and fairness evaluations for question answering (QA) systems such as the Bias Benchmark for QA (BBQ)Parrish et al. (2022) have been developed with Western, primarily English speaking contexts in mind. While these resources have been instrumental in revealing cultural and demographic biases, they do not adequately capture the unique social divisions, linguistic nuances, and historical power dynamics present in other regions. As a result, models deployed in low‐resource or non‐Western settings can exhibit untested and potentially more severe biases toward locally salient groups, such as caste, sect, or clan affiliations as supported in the following works Khandelwal et al. (2023), Ferrara (2023).
@@
Although there have been attempts to contextualize the original BBQ dataset in relation to the local context, little to no work has been done on a QA dataset tailored to the Pakistani context. KO-BBQ Jin et al. (2024) is a culturally adapted Korean version of the BBQ dataset, and is rooted in Korean culture, while its Chinese adapted version CBBQ Huang and Xiong (2024) captures nuances embedded within the Chinese culture. These datasets are not transferrable to Pakistani contexts, specifically due to the diverse social, cultural and language landscape of Pakistan. Pakistan’s rich cultural diversity stems from its multi-ethnic population spread across different provinces, each with distinct languages, traditions, views, and socio-economic status as shown in several studies Shah and Amjad (2011). With regional languages like Punjabi, Sindhi, Pashto, Balochi, and others coexisting alongside Urdu as the national language, and with deep rooted regional identities and cross regional biases, a one-size-fits-all dataset fails to capture the nuanced realities of the country. The effect is further propagated by Pakistan’s religious landscape, dominated by Islam which is split among sects like Barelvi, Deobandi, Christians, Hindus, and other minorities. This sectarian and interfaith diversity creates varied social norms and complicates uniform data representation. As Yaqin (2022) mentions, Urdu encodes social hierarchies through pronouns, honorifics, and register choices (Persian-Arabic vs. Hindustani vocabulary), signaling respect, status, and group membership. Gendered verb and adjective agreement further embeds masculine authority and feminine marginality. Modeling these formality markers in PakBBQ reveals how LLMs may reproduce structural biases along urban–rural, educational, and gender lines unique to Pakistan.
@@
To bridge this gap, we introduce PakBBQ (Pakistani Bias Benchmark for Question Answering), a culturally and regionally adapted bias benchmark for QA tailored to the Pakistani context. Building upon the original BBQ dataset, we imply a methodology similar to the one in KOBBQ to contextualize and adapt the dataset to Pakistani norms and culture. Templates were categorized into: Target Modified (TM) templates adapted for Pakistani context (e.g., replacing Western names with local counterparts), Sample Removed (SR) templates inapplicable locally, Directly Translated (DT) templates applicable locally and Newly Added (NA) templates capturing Pakistan specific biases (caste, sect, clan, regional affiliations), validated by native speakers. We also remove template categories irrelevant to Pakistani context, and add new categories such as Regional and Language Formality Biases, identified through large scale scrapping of various media articles, research papers, social blogs.
@@
  • PakBBQ Dataset: A collection of 214 templates instantiated into 17180 English and Urdu scripted QA pairs covering 8 bias dimensions specific to Pakistan.
@@
By releasing PakBBQ, covering eight bias dimensions (Age, Disability Status, Language Formality, Gender Identity, Physical Appearance, Regional, Religion, and Socioeconomic Status(SES)), we aim to enable more rigorous auditing and mitigation of social biases in QA models deployed in Pakistan, and to provide a blueprint for culturally sensitive bias benchmarks in other underrepresented regions. The dataset and code are available at out Github repository PakBBQ.
@@ 2.1 Bias Benchmarks in QA and Cross-Cultural Adaptations
Natural language processing models have been shown to inherit and even amplify societal biases present in their training data, which can manifest in question answering (QA) tasks as stereotypical or discriminatory outputs. Parrish et al. Parrish et al. (2022) introduced the Bias Benchmark for QA (BBQ) to evaluate such biases across nine social dimensions in U.S. English under both under informative and fully informative contexts. Subsequent frameworks, such as UnQover Li et al. (2020), employ underspecified questions to surface biases like gendered name–occupation associations, while pronoun based methods Zhao et al. (2018) reveal gender bias via pronoun usage, though these are less applicable in Urdu, which conveys gender through verb and adjective agreement rather than explicit pronouns.
@@ 3 Dataset Construction
3.1 Adaptation Strategy: DT, TM, NA and SR Categories
@@
Directly Translated (DT): Items in this category were translated into Urdu without significant changes, as their social and cultural contexts were already applicable to Pakistani society. These include examples with globally common scenarios like age-based assumptions or gender stereotypes.
@@
Newly Added (NA): This category includes examples specifically constructed for the Pakistani sociocultural landscape. These involve biases unique to Pakistan, such as sectarian affiliations, regional or ethnic identities (e.g., Sindhi, Baloch), and minority religious groups (e.g., Ahmadis, Hindus). We also incorporate formality biases specific to the Urdu language, since direct English prompts are often inadequate due to the complexity of Urdu’s morphological structure, gendered pronouns, and levels of formality in verbs, we instead use Roman Urdu to evaluate English responses. These examples aim to capture context-specific stereotypes not present in the original BBQ dataset.
@@ 3.2 Template Annotation
For the Newly Added (NA) templates, we employed a structured annotation process involving multiple annotators (undergraduate Pakistani university students), recruited as volunteers with native fluency in Urdu and English, and representing diverse regional backgrounds across Pakistan, to ensure both cultural relevance and consistency in identifying bias. Annotators were first briefed on the aims of the study and made explicitly aware of the potential risks of exposure to stereotypes, sexism, and other harmful biases contained within the templates. Each template was independently reviewed by each annotator in an isolated setting to prevent any external influences
@@ 3.3 Translation
For the translation of our dataset, we essentially translated the context, question and answer choices. Three translation models were evaluated, Facebook’s seamlessM4T-v2-large Barrault et al. (2023), Gemma3 27B  [^3] https://huggingface.co/google/gemma-3-27b-it Team et al. (2025) and Google Translate API  [^4] https://cloud.google.com/translate. Facebook and Gemma3 models frequently produced semantically distorted outputs, for example incorrectly translating “dark-skinned” as "andhera (i.e., ’darkness’) instead of Kaala (i.e., ’dark-skinned’)", and failed to preserve contextual cues across Q/As. After extensive manual comparison and evaluation, Google Translate API produced the most accurate and contextually sound translations, thus was chosen to translate our English version of the dataset to Urdu .
@@ 4 Experimental Setup
In the following section, we describe the inference methodology applied to evaluate the dataset across both languages an the detailed model selection to ensure model diversity.
@@ 4.2 Model Selection
We selected a diverse set of latest multilingual LLM models, capable of handling zero-shot question answering. We also ensures that a diversity was maintained in terms of model sizes as well, and the final models used were, DeepSeek-V3(671B)  [^5] https://deepseekv3.org , GPT-4.1-Nano  [^6] https://platform.openai.com/docs/models/gpt-4.1-nano , GPT-4.1-Mini [^7] https://platform.openai.com/docs/models/gpt-4.1-mini, GPT-4.1888https://platform.openai.com/docs/models/gpt-4.1, Gemini-2.0-Flash  [^9] https://deepmind.google/technologies/gemini/flash/ and Gemini-2.0-Flash-Lite [^10] https://deepmind.google/technologies/gemini/flash-lite/. While the exact parameters of some of these models have not been disclosed, we ensured that our evaluation covered a representative range of model scales, ranging from lightweight, midsized and large scale LLMs.
@@ 6.1 Accuracy Comparison
ENG GPT-4.1-Nano 0.80 0.83 0.68 0.81
GPT-4.1-Mini 0.82 0.87 0.62 0.87
GPT-4.1 0.82 0.88 0.49 0.89
DeepSeek-v3 0.85 0.91 0.55 0.92
gemini-2.0-flash-lite 0.88 0.93 0.61 0.94
gemini-2.0-flash 0.84 0.90 0.57 0.87
UR GPT-4.1-Nano 0.72 0.73 0.67 0.74
GPT-4.1-Mini 0.75 0.78 0.61 0.81
GPT-4.1 0.75 0.78 0.62 0.81
DeepSeek-v3 0.67 0.67 0.50 0.85
gemini-2.0-flash-lite 0.69 0.71 0.51 0.77
gemini-2.0-flash 0.81 0.85 0.61 0.88
Table 1: Accuracy comparison of LLMs across template types in English (ENG) and Urdu (UR)
@@ 6.2 Ambig vs Disambig Accuracy
All models gain substantially from disambiguation. Accuracy jumps by about 12 percentage points on average. Urdu models show higher variance than English ($\sigma\approx 0.11$ vs. 0.07), highlighting their sensitivity to prompt clarity. The results on a whole, underscore the value of explicit disambiguation, particularly in lower resource languages.
6.3 Negative vs Non-Negative Accuracy
@@
ENG GPT-4.1-Nano 0.83 0.78
GPT-4.1-Mini 0.83 0.82
GPT-4.1 0.81 0.81
DeepSeek-v3 0.85 0.84
Gemini-2.0-Flash-Lite 0.89 0.87
Gemini-2.0-Flash 0.85 0.82
UR GPT-4.1-Nano 0.73 0.71
GPT-4.1-Mini 0.77 0.74
GPT-4.1 0.77 0.74
DeepSeek-v3 0.69 0.64
Gemini-2.0-Flash-Lite 0.71 0.66
Gemini-2.0-Flash 0.81 0.81
Table 2: Performance Comparison of LLMs on Negative and Non-Negative Questions across Urdu(UR) and English(ENG
@@ 6.4 Bias Scores Across Categories
Accuracy is not enough to evaluate biases in LLMs as we must also evaluate how biased or counter-biased the model is if it does not choose the Unknown option. We apply the bias score metric used in BBQParrish et al. (2022). The provided heatmaps in Figure  6 illustrate bias scores for models evaluated on both English and Urdu across several bias categories, including Age, Disability Status, Gender Identity, Physical Appearance, Regional Bias, Religion, and Socioeconomic Status (SES), under both ambiguous and disambiguated contexts. The bias score (1-1 to $1$) measures a model’s preference between biased and counter-biased choices (excluding “Unknown”): a score of 0 indicates no preference, values near $1$ indicate bias, and values near 1-1 indicate counter-bias.
@@ 7.4 Disambiguation as a Bias Mitigation Strategy
Explicit disambiguation yields an average accuracy gain of 12 pp across all models (e.g., GPT4.1-Mini in Urdu rises from 64 % to 87 %). This suggests that much of the bias in ambiguous prompts stems from models’ reliance on learned probabilistic defaults rather than deliberate reasoning, and that clear contextual cues enable them to override these assumptions. In practice, disambiguation offers a straightforward prompt‐engineering technique, especially valuable for lower‐resource languages like Urdu, though it may be less feasible when explicit information is unavailable. The stronger effect observed in Urdu also points to language‐specific bias mechanisms, warranting further investigation into how linguistic structure and culture shape model behavior.
@@ 8 Conclusion
This paper introduces PakBBQ, the first culturally adapted bias benchmark for evaluating large language models in the Pakistani context. Our findings reveal substantial performance disparities between English and Urdu (7-17 percentage points accuracy gap) and demonstrate that bias evaluation cannot rely on simple translation of Western frameworks. Pakistan specific bias categories showed consistently poor performance, highlighting the necessity of culturally grounded evaluation. However, we identify practical mitigation strategies, explicit disambiguation improves accuracy by 12pp on average, while negative question framing reduces stereotypical responses—both effects being stronger in Urdu than English. These findings challenge the transferability assumption in AI bias evaluation and provide immediate prompt engineering solutions for more equitable multilingual AI deployment. Our work establishes a replicable methodology for developing culturally specific bias benchmarks, emphasizing the urgent need to move beyond English centric evaluation frameworks toward inclusive approaches that address the diverse realities of global AI systems and paves the way for similar adaptations in other South Asian contexts.
@@
While PakBBQ represents a first step towards culturally grounded bias evalution for Pakistani QA systems, several limitation should be noted. Our dataset is primarily limited to Urdu and English scripts and does not cater towards major regional languages (eg., Punjabi, Sindhi, Pashto, Balochi), which may exhibit distinct bias patterns. Secondly our evaluation adopts a zero‐shot prompting strategy with a fixed “You are a Pakistani person” system prompt and relies on the assumption that all models interpret formality and honorific cues consistently in both languages; in practice, morphological and register mismatches may introduce noise. Furthermore, errors and context misalignment is possible during the translation of the dataset, translation errors can arise especially for context sensitive terms like skin tone or sectarian identifiers which may affect bias measurements downstream.
@@
Future Work
Building on the limitations outlined in this work, future research may extend PakBBQ to additional regional languages spoken in Pakistan, such as Pashto, Sindhi, Punjabi, and Balochi, which collectively represent a significant portion of the population. Additionally, evaluating LLMs in multimodal or cross-modal settings presents a promising avenue to assess biases beyond text based QA; for instance, integrating such benchmarks with VLMs could probe stereotypical representations in generated images or media, while audio evaluations might examine biases in speech related tasks, including accent or formality variations. Moreover, future work could also explore chain-of-thought prompting to compare reasoning paths across languages, examining how differences in training data, linguistic structures, or cultural priors affect model responses and bias mitigation on such datasets.